how much gallons is 200 pints? show your work.
Answer by Tobiasz(54)

Ok
200 pints into gallons : ] 1+US+gallon+=+8+US+pints
1 gallons equals 8 pints
all you do is 200 pints divided by 8 and that equals to 25
______________________________________________________________
ok lets do this one 540 pints into gallons
540 / 8 = 67.5
or this one:
______________________________________________
2000 pints into gallons
2000/8 = 250
HOPE I HELPED =p HERE IS A CHART TO HELP! CYA!
http://www.lemelange.com/conversion_chart.htm <----- the chart